OpenStudy (anonymous):

8th grade math help???? 5 questions???? Medals!!!!!!!!!!!!!!!

OpenStudy (anonymous):

\[4(2x+3)=2(3x-4)\]

rishavraj (rishavraj):

8x + 12 = 6x - 8 2x = -20 x = -10